# Evaluating the Defense Potential of Machine Unlearning against Membership Inference Attacks

### v3

@@ Abstract
Membership Inference Attacks (MIAs) pose a significant privacy risk, as they enable adversaries to determine whether a specific data point was included in the training dataset of a model. While Machine Unlearning is primarily designed as a privacy mechanism to efficiently remove private data from a machine learning model without the need for full retraining, its impact on the susceptibility of models to MIA remains an open question. In this study, we systematically assess the vulnerability of models to MIA after applying state-of-art Machine Unlearning algorithms. Our analysis spans four diverse datasets (two from the image domain and two in tabular format), exploring how different unlearning approaches influence the exposure of models to membership inference. The findings highlight that while Machine Unlearning is not inherently a countermeasure against MIA, the unlearning algorithm and data characteristics can significantly affect a model’s vulnerability. This work provides essential insights into the interplay between Machine Unlearning and MIAs, offering guidance for the design of privacy-preserving machine learning systems.
@@ 1 Introduction
As machine learning (ML) models continue to advance, their applications have become widespread across various industries, such as healthcare nikolaidis2023federated , telecommunications tsiolakis2024carbon , biological sciences pavlidis2023extensive and beyond. These models have revolutionized how we approach tasks, making decision-making processes more efficient and accurate. However, the extensive use of ML raises concerns about the reliance on large volumes of user data, often containing sensitive or personal information. With the increasing number of privacy incidents where unauthorized parties gain access to this sensitive data, the need for robust privacy-preserving mechanisms has become critical.
One of the primary privacy vulnerabilities of ML models is defending against Membership Inference Attacks (MIAs) shokri2017membership . In such attacks, adversaries aim to determine whether a specific data point was part of a model’s training dataset. This information can potentially lead to the leakage of sensitive data, posing a significant threat to user privacy, especially when sensitive data such as medical records are involved. For instance, if a healthcare model can be queried to confirm whether specific patient records were used during its training, it could expose confidential health information about individuals. The implications of such attacks extend to other domains such as social media, finance and security. Therefore, understanding and mitigating the risks posed by MIAs is crucial for ensuring the privacy and safety of users.
In recent years, Machine Unlearning Cao has gained attention as a potential method for addressing privacy concerns in ML applications. Machine Unlearning allows service providers to erase specific data from a trained model without needing to retrain the model from scratch. This technique is particularly useful in the context of privacy regulations like the General Data Protection Regulation (GDPR), which requires the deletion of user data upon request. By using such erasing algorithms, organizations can efficiently comply with such legal requirements while minimizing computational costs and time associated with full model retraining. Furthermore, Machine Unlearning can be utilized as a security measure zhang2023review , since it reduces the influence of deleted data points on model predictions, potentially lowering the model’s susceptibility to adversarial attacks like MIAs.
@@
Despite its appeal, the relationship between Machine Unlearning and a model’s vulnerability to MIAs remains largely unexplored. While Machine Unlearning primarily focuses on the efficient removal of data to protect user privacy, it is unclear how this process influences the model’s susceptibility to MIAs liu2024threats . This knowledge gap is crucial because even after certain data has been ”unlearned”, models may still retain enough information about the deleted samples to be vulnerable to MIAs. As such, merely applying unlearning techniques may not suffice to mitigate these risks unless the process is carefully evaluated and tailored to address the specific threats posed by MIAs.
@@
This work aims to fill this gap by assessing the impact of Machine Unlearning on a model’s vulnerability against MIAs. We evaluate the efficacy of state-of-the-art Machine Unlearning algorithms across four distinct datasets, each representing different applications. Our analysis focuses on how unlearning methods affect a model’s exposure to MIAs and whether certain unlearning strategies are more effective than others in reducing this risk. Our findings suggest several important insights. First, while Machine Unlearning can effectively remove targeted data, it does not inherently serve as a defense mechanism against MIAs. In some cases, the removal of data via unlearning may even introduce subtle vulnerabilities that could increase the model’s susceptibility to adversarial attacks. Second, the effectiveness of unlearning techniques in reducing vulnerability to MIAs varies significantly depending on the unlearning algorithm used and the use-case scenario. Hence, designing a robust Machine Unlearning solution requires careful consideration of the the specific nature of the data being unlearned.
@@ 2.1 Membership Inference Attacks (MIAs)
MIAs present a significant privacy challenge in ML models. Specifically, a MIA involves an adversary trying to determine if a specific data point was part of the model’s training dataset hu2021membership . Given access to the model’s outputs, the adversary can exploit differences in the model’s behavior when given data that was part of the training set versus data that was not (see Fig. 1). This capability enables the adversary to deduce whether specific instances are in the training set, potentially revealing sensitive information about individuals.
@@
The success of a MIA depends on how effectively the adversary can differentiate between members of $\mathcal{D}_{train}$ and non-members based on the model’s predictions. Such attacks pose significant privacy risks as it could potentially leak sensitive information, especially in models trained on personal or confidential data, such as medical or financial records. In this work, we implement a white-box attack, where the adversary has knowledge of the target model’s architecture and the distribution of the dataset niu2024survey . This type of attack is more powerful, as it exploits deeper insights into the model, making it easier to infer membership.
@@ 2.2 Machine Unlearning
Machine Unlearning refers to the efficient process of removing the influence of specific data points from a trained ML model. This concept is particularly important in situations where users request the deletion of their data, especially under privacy regulations like the General Data Protection Regulation (GDPR) 10488864 . The challenge is to ensure that, after the data is removed, the model no longer retains any trace of the deleted data, while avoiding the expensive process of retraining the entire model from scratch (see Fig. 2).
@@ 2.3 Influence of Machine Unlearning on Membership Inference Attacks (MIAs)
While Machine Unlearning is designed to remove data points from a model’s training set, its influence on the model’s vulnerability to MIA is not straightforward. The process of unlearning aims to erase the impact of the data to be forgotten ($\mathcal{D}_{forget}$), but this can either reduce or exacerbate the model’s susceptibility to MIAs depending on how the unlearning is implemented Chen .
@@
3 Related Work
3.1 MIA
Shokri et al. shokri2017membership introduced the concept of MIAs, demonstrating that adversaries could determine whether a particular data point was part of the training dataset by observing the model’s outputs. Their work laid the groundwork for subsequent research in this domain.
Salem et al. salem2018ml extended the MIA concept to more practical scenarios, showing that even with limited knowledge of the target model, adversaries could perform effective MIA. They explored different attack models, including black-box and white-box scenarios, highlighting the pervasive risk of MIA.
A comprehensive survey by Hu et al. hu2021membership details the variety of MIA techniques across different ML models, including classification and generative models. Their work also outlines the defences proposed to mitigate these attacks, emphasizing the ongoing need for robust privacy-preserving mechanisms. Other studies, such as Zhong2022Understanding explore the privacy unfairness in MIA and how certain defence mechanisms, like differential privacy and dropout, affect different subgroup data.
Further advancements in MIA, such as the MIA-BAD framework proposed by Banerjee et al banerjee2023miabad leverage federated learning as a potential defence mechanism against batch-wise MIA attacks. Similarly, Shi et al. shi2023scalemia introduce Scale-MIA, a scalable attack capable of recovering training data even in secure federated learning environments, demonstrating the persistent risks to user privacy.
3.2 Machine Unlearning
Cao et al. Cao introduced the foundational concept of machine unlearning, where they focused on the efficient removal of data from ML models. Their method transforms models like Naive Bayes into a summation-based structure, enabling the deletion of specific data points without requiring full retraining of the entire model.
Building on this foundational work, Bourtoule et al. Bourtoule introduced a more scalable unlearning framework called SISA in 2019. The core idea behind SISA training is to divide the training dataset into several disjoint shards and train sub-models on these separate shards. This approach enhances scalability since, when a data point needs to be removed, only the sub-models trained on the affected shards need retraining.
Ginart et al. Ginart tackled unlearning in the context of unsupervised learning, specifically k-means clustering. Their method quantizes data points into discrete clusters, enabling efficient unlearning by allowing modifications or deletions of the clusters containing the data points in question, without needing to retrain the entire model.
While these approaches focus on unlearning efficiency, it is equally important to minimize the negative impact on model accuracy. Addressing this concern, Perifanis et al. perifanis2024sftc proposed SFTC, an approach that fine-tunes the model on the remaining instances while intentionally confusing the model about the forget set, thus balancing unlearning with preserving predictive performance.
Moving to the context of logit-based classifiers, Baumhauer et al. Baumhauer explored machine unlearning by leveraging the structure of logit-based models. This approach provides an efficient method for removing data points, avoiding full retraining while ensuring data removal.
To tie all these efforts together, Xu et al. Xu conducted a comprehensive survey of various machine unlearning techniques. Their survey covers the broad range of methods introduced so far, analyzing the trade-offs between accuracy, efficiency, and scalability, and highlighting the ongoing challenges in this evolving field.
3.3 Machine Unlearning and MIA
Machine unlearning has emerged as a promising defense mechanism against MIAs, where adversaries attempt to infer whether specific data points were part of the training set. Several studies have explored various aspects of unlearning in this context, shedding light on both its potential and its limitations.
For instance, Sommer et al. Sommer investigate probabilistic verification methods to defend against MIA in unlearning systems, emphasizing the need for strong verification protocols to prevent data leakage. Their work underscores the necessity of robust mechanisms to ensure the integrity of unlearning processes.
Also, Doughan and Itani DoughanItani provide a comparative analysis of various machine unlearning methods, focusing on their effectiveness in hindering MIAs. They emphasize the importance of future research in this area, particularly in the context of different models and datasets. In their study, they examine machine unlearning techniques applied to Resnet-18 models trained with the CIFAR-10 dataset. By comparing Fine-Tuning and Fisher Noise-based Impair-Repair methods, they demonstrate that the latter significantly reduces MIA scores while maintaining model accuracy.
However, even with such advancements, new challenges continue to emerge. Lu et al. Lu introduce label-only MIAs, highlighting a novel threat to machine unlearning systems. These attacks enable adversaries to infer membership without relying on posterior probabilities, thus complicating existing defense strategies.
In addition to these concerns, Chen et al. Chen delve deeper into the privacy risks associated with unlearning. They show that MIAs can still exploit unlearning processes, even when sensitive data is expected to be removed. The authors propose a novel MIA that targets models post-unlearning by leveraging differences in model outputs before and after the unlearning process, demonstrating how attackers can infer which data points were likely part of the original training set.
Finally, Niu et al. NIU2024404 offer a comprehensive review of the landscape of MIAs. Surveying over 200 publications, they systematically examine various defense strategies, including machine unlearning, to mitigate these attacks. Their findings highlight the importance of careful evaluation in the unlearning process, noting that incomplete or ineffective unlearning can leave models vulnerable to residual traces of removed data.

### v4

@@ Abstract
Membership Inference Attacks (MIAs) pose a significant privacy risk by enabling adversaries to determine if a specific data point was part of a model’s training set. This work empirically investigates whether Machine Unlearning (MU) algorithms can function as a targeted, active defense mechanism, in scenarios where a privacy audit identifies specific classes or individuals as highly susceptible to MIAs post-training. By ’dulling’ the model’s categorical memory of these samples, the process effectively mitigates the membership signal and reduces the MIA success rate for the most vulnerable users. We evaluate the defense potential of three MU algorithms, Negative Gradient (neg grad), SCalable Remembering and Unlearning unBound (SCRUB), and Selective Fine-tuning and Targeted Confusion (SFTC), across four diverse datasets and three complexity-based model groups. Our findings reveal that MU can function as a countermeasure against MIAs, though its success is critically contingent on algorithm choice, model capacity, and a profound sensitivity to learning rates. While Negative Gradient often induces a generalized degradation of membership signals across both forget and retain set, SFTC identifies a critical “divergence effect” where targeted forgetting reinforces the membership signal of retained data. Conversely, SCRUB provides a more balanced defense with minimal collateral impact on MIA perspective.
@@ 1 Introduction
As machine learning (ML) models continue to advance, their applications have become widespread across various industries, such as healthcare Nikolaidis and Efraimidis (2024), telecommunications Tsiolakis et al. (2024), biological sciences Pavlidis et al. (2023) and beyond. These models have revolutionized how we approach tasks, making decision-making processes more efficient and accurate. However, the extensive use of ML raises concerns about the reliance on large volumes of user data, often containing sensitive or personal information. With the increasing number of privacy incidents where unauthorized parties gain access to this sensitive data, the need for robust privacy-preserving mechanisms has become critical.
One of the primary privacy vulnerabilities of ML models is defending against MIAs Shokri et al. (2017). In such attacks, adversaries aim to determine whether a specific data point or individual client was part of a model’s training dataset Nikolaidis and Efraimidis (2025). For instance, if a social media model can be queried to confirm whether specific users’ image were used during its training, it could expose confidential information about individuals. Therefore, understanding and mitigating the risks posed by MIAs is crucial for ensuring the privacy and safety of users. Beyond mere compliance, this study investigates  [^1] Code available at: https://github.com/TheoTsio/MIA_vs_Unlearning.git whether MU  Cao and Yang (2015) can be repurposed as an active, targeted defense against such vulnerabilities. While both domains center on the residual influence of specific data points, MU traditionally allows service providers to erase data without retraining from scratch. While primarily utilized for regulatory compliance, such as fulfilling the “right to be forgotten” imposed by the General Data Protection Regulation (GDPR), MU has the potential to function as a latent security measure Zhang et al. (2023). By strategically applying unlearning to “patch” high-risk data points, we can evaluate its efficacy in masking the traces of training membership and hardening the model against adversarial inference.
@@
Despite its appeal, the relationship between Machine Unlearning and a model’s vulnerability to MIAs remains largely unexplored, and these kind of attacks mostly used as an evaluation metric of the MU quality Tu et al. (2025). While MU primarily focuses on the efficient removal of data to protect user privacy, it is unclear how this process influences the model’s susceptibility to MIAs Liu et al. (2024). This knowledge gap is crucial because even after certain data has been “unlearned”, models may still retain enough information about the deleted samples to be vulnerable to MIAs. As such, merely applying unlearning techniques may not suffice to mitigate these risks unless the process is carefully evaluated and tailored to address the specific threats posed by MIAs.
@@
This work aims to contribte to this research gap by assessing the impact of Machine Unlearning on a model’s vulnerability against MIAs. We evaluate the efficacy of state-of-the-art Machine Unlearning algorithms across four distinct datasets, each representing different applications. Our analysis focuses on how unlearning methods affect a model’s exposure to MIAs and whether certain unlearning strategies are more effective than others in reducing this risk. Our findings suggest several important insights. First, while Machine Unlearning can effectively remove targeted data, it does not inherently serve as a defense mechanism against MIAs. In some cases, the removal of data via unlearning may even introduce subtle vulnerabilities that could increase the model’s susceptibility to adversarial attacks. Second, the effectiveness of unlearning techniques in reducing vulnerability to MIAs varies significantly depending on the unlearning algorithm used and the use-case scenario. Hence, designing a robust Machine Unlearning solution requires careful consideration of the the specific nature of the data being unlearned.
@@
2 Key Takeaways of the paper
  • $\bullet$
    Defense Focus: Explicitly frames Machine Unlearning (MU) as an active countermeasure against privacy threats that are recognized in the post-trained model.
  • $\bullet$
    Comparative Analysis: Highlights the distinct performance profiles of SCRUB and SFTC, identifying a trade-off between the stability of the former and the high selectivity of the latter.
  • $\bullet$
    Experimental Scope: Validates findings across a broad experimental matrix comprising four diverse datasets (image and tabular) and three distinct model complexity groups.
  • $\bullet$
    Continuous Auditing Requirement: Establishes that evaluating MIA accuracy at every epoch is essential to capture unlearning dynamics and identify the optimal “stopping point” before “over-unlearning” degrades model utility.
  • $\bullet$
    MIA Accuracy Trade-offs: Reveals that a successful reduction in MIA Forget Accuracy can paradoxically lead to an increase in MIA Retain Accuracy. While the model becomes less vulnerable regarding the forgotten set, intensive fine-tuning can reinforce the membership signal of the retain set, potentially compromising its privacy.
3 Related Work
3.1 MIA
Shokri et al. Shokri et al. (2017) introduced the concept of MIAs, demonstrating that adversaries could determine whether a particular data point was part of the training dataset by observing the model’s outputs. Their work laid the groundwork for subsequent research in this domain.
Salem et al. Salem et al. (2018) extended the MIA concept to more practical scenarios, showing that even with limited knowledge of the target model, adversaries could perform effective MIA. They explored different attack models, including black-box and white-box scenarios, highlighting the pervasive risk of MIA.
A comprehensive survey by Hu et al. Hu et al. (2021) details the variety of MIA techniques across different ML models, including classification and generative models. Their work also outlines the defences proposed to mitigate these attacks, emphasizing the ongoing need for robust privacy-preserving mechanisms. Other studies, such as Zhong et al. (2022) explore the privacy unfairness in MIA and how certain defence mechanisms, like differential privacy and dropout, affect different subgroup data.
Further advancements in MIA, such as the MIA-BAD framework proposed by Banerjee et al Banerjee et al. (2023) leverage federated learning as a potential defence mechanism against batch-wise MIAs. Similarly, Shi et al. Shi et al. (2023) introduce Scale-MIA, a scalable attack capable of recovering training data even in secure federated learning environments, demonstrating the persistent risks to user privacy.
3.2 Machine Unlearning
Cao et al. Cao and Yang (2015) introduced the foundational concept of machine unlearning, where they focused on the efficient removal of data from ML models. Their method transforms models like Naive Bayes into a summation-based structure, enabling the deletion of specific data points without requiring full retraining of the entire model.
Building on this foundational work, Bourtoule et al. Bourtoule et al. (2021) introduced a more scalable unlearning framework called SISA in 2019. The core idea behind SISA training is to divide the training dataset into several disjoint shards and train sub-models on these separate shards.
Ginart et al. Ginart et al. (2019) tackled unlearning in the context of unsupervised learning, specifically k-means clustering. Their method quantizes data points into discrete clusters, enabling efficient unlearning by allowing modifications or deletions of the clusters containing the data points in question, without needing to retrain the entire model.
While these approaches focus on unlearning efficiency, it is equally important to minimize the negative impact on model accuracy. Addressing this concern, Perifanis et al. Perifanis et al. (2024) proposed SFTC, an approach that fine-tunes the model on the remaining instances while intentionally confusing the model about the forget set, thus balancing unlearning with preserving predictive performance.
Moving to the context of logit-based classifiers, Baumhauer et al. Baumhauer et al. (2022) explored machine unlearning by leveraging the structure of logit-based models. This approach provides an efficient method for removing data points, avoiding full retraining while ensuring data removal.
To tie all these efforts together, Xu et al. Xu et al. (2023) conducted a comprehensive survey of various machine unlearning techniques. Their survey covers the broad range of methods introduced so far, analyzing the trade-offs between accuracy, efficiency, and scalability, and highlighting the ongoing challenges in this evolving field.
3.3 Machine Unlearning and MIA
Several studies have explored various aspects of MU with different ideas of adversarial attacks and more specific to MIAs, shedding light on both its potential and its limitations.
For instance, Sommer et al. Sommer et al. (2022) investigate probabilistic verification methods to defend against MIA in unlearning systems, emphasizing the need for strong verification protocols to prevent data leakage. Their work underscores the necessity of robust mechanisms to ensure the integrity of unlearning processes.
Also, Doughan and Itani Doughan and Itani (2024) provide a comparative analysis of various machine unlearning methods, focusing on their effectiveness in hindering MIAs. They emphasize the importance of future research in this area, particularly in the context of different models and datasets. In their study, they examine machine unlearning techniques applied to Resnet-18 models trained with the CIFAR-10 dataset. By comparing Fine-Tuning and Fisher Noise-based Impair-Repair methods, they demonstrate that the latter significantly reduces MIA scores while maintaining model accuracy.
However, even with such advancements, new challenges continue to emerge. Lu et al. Lu et al. (2022) introduce label-only MIAs, highlighting a novel threat to machine unlearning systems. These attacks enable adversaries to infer membership without relying on posterior probabilities, thus complicating existing defense strategies.
In addition to these concerns, Chen et al. Chen et al. (2021) delve deeper into the privacy risks associated with unlearning. They show that MIAs can still exploit unlearning processes, even when sensitive data is expected to be removed. The authors propose a novel MIA that targets models post-unlearning by leveraging differences in model outputs before and after the unlearning process, demonstrating how attackers can infer which data points were likely part of the original training set.
Finally, Niu et al. Niu et al. (2024a) offer a comprehensive review of the landscape of MIAs. Surveying over 200 publications, they systematically examine various defense strategies, including machine unlearning, to mitigate these attacks. Their findings highlight the importance of careful evaluation in the unlearning process, noting that incomplete or ineffective unlearning can leave models vulnerable to residual traces of removed data.
@@ 4.1 Membership Inference Attacks (MIAs)
MIAs present a significant privacy challenge in ML models. Specifically, a MIA involves an adversary trying to determine if a specific data point was part of the model’s training dataset Hu et al. (2021). Given access to the model’s outputs, the adversary can exploit differences in the model’s behavior when given data that was part of the training set versus data that was not (see Fig. 1). This capability enables the adversary to deduce whether specific instances are in the training set, potentially revealing sensitive information about individuals.
@@
The success of a MIA depends on how effectively the adversary can differentiate between members of $\mathcal{D}_{train}$ and non-members based on the model’s predictions. Such attacks pose significant privacy risks as it could potentially leak sensitive information, especially in models trained on personal or confidential data, such as medical or financial records. In this work, we implement a white-box attack, where the adversary has knowledge of the target model’s architecture and the distribution of the dataset Niu et al. (2024b). This type of attack is more powerful, as it exploits deeper insights into the model, making it easier to infer membership.
4.2 Shadow Models
To execute MIAs, we employ shadow models to simulate the target model’s behavior and establish a ground truth for membership. The adversary trains multiple proxy models on datasets with known membership, allowing them to observe the distinct output probability distributions, often characterized by higher confidence or lower entropy, associated with training data versus unseen data. These observations are used to train an attack classifier that learns to recognize the “membership signal”. In the context of machine unlearning, shadow models are critical for auditing; they allow for a direct comparison between a model that has “unlearned” a specific data point and a “retrained-from-scratch model. If an attack model can still distinguish a forgotten sample from a truly unseen one, it indicates a privacy residual, suggesting that the unlearning algorithm has failed to completely erase the influence of the deleted data.
@@ 4.3 Machine Unlearning
The operational workflow of Machine Unlearning (as illustrated in Fig. 2) involves identifying a target subset of data, defined here as a specific percentage of the original training set, to be removed from a pre-trained model. Once the target data is selected, the unlearning algorithm is applied to adjust the model’s parameters. A critical challenge throughout this procedure is the preservation of the model’s utility; if the predictive capacity is substantially degraded following the unlearning process, the procedure loses its practical and functional significance. Therefore, an effective unlearning execution must achieve a state where specific data influence is erased without inducing a total predictive collapse.
@@ 4.4 Influence of Machine Unlearning on Membership Inference Attacks (MIAs)
While Machine Unlearning is designed to remove data points from a model’s training set, its influence on the model’s vulnerability to MIA is not straightforward. The process of unlearning aims to erase the impact of the data to be forgotten ($\mathcal{D}_{forget}$), but this can either reduce or exacerbate the model’s susceptibility to MIAs depending on how the unlearning is implemented Chen et al. (2021).
@@
4.5 White box and Black box MIA
Membership Inference Attacks (MIA) are categorized based on the information available to the adversary regarding the target model’s parameters and architecture.
  • $\bullet$
    Black-box MIA: In this restricted setting, the adversary can only interact with the target model via an API, providing an input x and receiving the corresponding output y, which is typically a vector of prediction probabilities (posteriors). The attacker exploits the statistical differences in these confidence scores, as models often exhibit higher confidence and lower entropy for members of the training set compared to unseen data. This study utilizes shadow models to simulate this behavior and train an attack classifier without internal model access.
  • $\bullet$
    Grey-box MIA: This intermediate setting assumes the adversary possesses partial information about the target model or its training process. This may include knowledge of the model’s architecture (e.g., specific ResNet or MLP structures) or the underlying data distribution, without having direct access to the model’s weights. By aligning the shadow model’s architecture with that of the target, the attacker can more accurately replicate the target model’s decision boundaries, leading to a more effective inference than a strictly black-box approach.
  • $\bullet$
    White-box MIA: In a white-box scenario, the adversary has full access to the target model’s internal components, including its weights, specific architecture, and intermediate layer activations. This allows for more sophisticated attacks that exploit the model’s internal gradients or “memorization” patterns within specific neurons. White-box attacks generally provide an upper bound on privacy leakage as they utilize the maximum possible information available.
4.6 Unlearning Algorithms Families
Machine unlearning algorithms are generally classified into two broad categories based on their approach to removing the influence of specific data points from a trained model.
  • $\bullet$
    Exact Unlearning: This paradigm aims to produce a model that is statistically indistinguishable from one trained from scratch without the “forget set.” While retraining is the most straightforward example, more efficient exact methods (e.g., SISA) leverage specialized data partitioning and partial retraining to reduce computational costs. Because these methods provide a theoretical guarantee of data removal, they serve as the gold standard for privacy compliance.
  • $\bullet$
    Unlearning: This more common category seeks to efficiently approximate the ideal unlearned model through heuristics and optimization without full retraining. These methods typically involve targeted parameter updates to reduce the impact of the forgotten data to a practical level. Our study focuses on three representative approximate methods:
    • $\bullet$
      NegGrad (Negative Gradient): Serves as a fundamental gradient-based baseline. It is the standard approach for direct parameter manipulation and is essential for establishing a lower bound on unlearning performance.
    • $\bullet$
      SCRUB: Represents the state-of-the-art in teacher - student frameworks. It is widely cited Kurmanji et al. (2023) for its ability to balance the removal of “forget” information with the preservation of “retain” utility via knowledge distillation.
    • $\bullet$
      SFTC: A recent and promising approach that optimizes unlearning fidelity Perifanis et al. (2024). represents the “selective update” family of algorithms, which are currently favored for their efficiency and ability to handle complex data distributions.
4.7 Metrics
4.7.1 Forget Accuracy
Forget Accuracy is a metric used to evaluate the effectiveness of a machine unlearning algorithm. It is defined as the accuracy of the unlearned model when evaluated on the forget set, the specific subset of data that was targeted for removal. A successful unlearning process aims to minimize this accuracy, ideally reducing it to a level equivalent to that of a model that was never trained on this data in the first place (i.e., no better than random guessing or its baseline generalization performance on unseen data). A low Forget Accuracy indicates that the model no longer retains the specific patterns learned from the forgotten data.
4.7.2 Retain Accuracy
Retain Accuracy is a metric that measures the utility preservation of a machine unlearning algorithm. It is defined as the accuracy of the unlearned model when evaluated on the retain set—the portion of the original training data that was not targeted for removal. The primary goal is to maintain a Retain Accuracy that is as high as possible, and ideally, statistically indistinguishable from the accuracy of the original model on this same set. This metric ensures that the unlearning process, while removing specific information, does not cause “catastrophic forgetting” or degrade the model’s performance on the remaining, valid data.
4.7.3 MIA Accuracy
MIA Accuracy refers to the accuracy of a Membership Inference Attack (MIA). Meaning the number of samples that correctly identified as part of the training dataset. So if we want to quantify it, this metric is the number of samples that correctly identified as part or not of the original dataset compared to total number of samples queried.
$MIA\ Acc=\frac{Correct\ Predictions}{Total\ Samples}$ (2)
where Total Samples are all the samples we are testing, both members and non members.
4.7.4 MIA Forget Accuracy
MIA Forget Accuracy is a privacy-verification metric used to quantify the residual influence of deleted data by evaluating the success rate of a Membership Inference Attack on the forget set. It measures the proportion of samples within the forget set that the adversary can still correctly identify as having been part of the original training distribution. By applying the MIA specifically to the forget set, we can empirically determine the extent to which “deleted” samples continue to leave recognizable traces or unique signatures within the model’s parameters. A successful unlearning process aims to reduce this accuracy to a baseline level, confirming that the model’s behavior on the forgotten data is indistinguishable from its behavior on samples it has never encountered.
$MIA\ Accuracy_{forget}=\frac{C_{F}+C_{U}}{|D_{forget}|+|D_{unseen}|}$ (3)
where $C_{F}$ and $C_{U}$ represent correct predictions for forget and unseen samples, respectively. A successful unlearning process aims to reduce this accuracy to a baseline of 50 %, confirming that the model’s behavior on the forgotten data is indistinguishable from its behavior on samples it has never encountered.
4.7.5 MIA Retain Accuracy
MIA Retain Accuracy serves as a measure of how the unlearning process affects the membership privacy of the data intended to be kept. It is defined as the accuracy of an MIA when evaluated on the retain set—the samples that were not targeted for removal. This metric tracks how many samples in the retain set continue to be correctly identified as members of the training distribution after the model has been modified. A significant deviation in MIA Retain Accuracy compared to the original model’s baseline suggests that the unlearning process may have overly perturbed the model’s parameters, potentially reducing its confidence or changing its behavior on the valid remaining data in a way that alters its privacy profile.
$MIA\ Accuracy_{retain}=\frac{C_{R}+C_{U}}{|D_{retain}|+|D_{unseen}|}$ (4)
where $C_{R}$ and $C_{U}$ represent correct predictions for retain and unseen samples, respectively.